# Comparison of bar formation mechanisms II: does a tidally induced bar grow faster than an internally developed bar?

### v1

@@ Abstract
Bar structures can form internally due to the instability of their host galaxies or externally due to perturbations from other galaxies. We systematically quantify the growth timescales ($\tau_{\mathrm{bar}}$) of bars formed through these two mechanisms with a series of controlled $N$-body simulations. In galaxies susceptible to bar instability, tidally induced bars display $\tau_{\mathrm{bar}}$ values comparable to those of internally developed bars within the same disk. Tidal perturbations promote/delay bar formation by advancing/postponing its onset, but the growth rate of the bar structure remains largely unchanged. In these interaction scenarios, the bar formation is still driven primarily by the galaxy’s internal nature that remains unaffected by tidal perturbations. As the external perturbation wave reaches the galaxy’s center, it evokes a “seed bar” that is then swing amplified. In this scenario, the onset of bar formation is advanced. Conversely, bar formation may be delayed if the external perturbation wave is out of phase with the pre-existing spontaneously developed seed bar, which causes destructive interference and limits the bar growth. In the hot disk model that resists bar formation in isolation, $\tau_{\mathrm{bar}}$ of the tidally induced bar correlates with the strength of the perturbation. The bar growth in this model deviates from the exponential profile and is better described by a linear function. The strictly-speaking “tidally induced bars” may not adhere to the swing amplification mechanism that predicts an exponential bar growth. Their preference for linear growth contrasts with bars formed in galaxies inherently susceptible to bar instability.
@@ 1 Introduction
Bars are a common feature in spiral galaxies. Locally, about half of the disk galaxies possess bars in optical bands, with this proportion rising to two-thirds in infrared bands (Marinova & Jogee, 2007; Menéndez-Delmestre et al., 2007; Erwin, 2018; Lee et al., 2019). Although a smaller fraction of galaxies have bars at higher redshifts, they still represent 13% in the range $2<z\leq 3$ (Le Conte et al., 2024) and appear as early as $z\sim 4$ (Guo et al., 2024). Bars significantly influence the evolution of their host galaxies, including triggering gas inflow, affecting star formation, and contributing to the formation of pseudo-bulges (Masters et al., 2011; Li et al., 2015; Lin et al., 2017, 2020; Iles et al., 2022). The Milky Way also has a bar (e.g. de Vaucouleurs, 1964; Blitz & Spergel, 1991), and its boxy-bulge and gas dynamics are directly linked to this bar (e.g. Shen et al., 2010; Li et al., 2018). The prevalence and impact of bars on galaxies highlight the importance of studying their formation and evolution.
@@
Galactic bars can form via internal or external mechanisms. Internally, bars can develop spontaneously due to the gravitational instability of the disks (Hohl, 1971; Ostriker & Peebles, 1973; Sellwood, 2014; Łokas, 2019a). A widely recognized mechanism for spontaneous bar formation is the swing amplification loop, initially developed by Goldreich & Lynden-Bell (1965); Julian & Toomre (1966) and further elaborated by Toomre (1981). Binney (2020) revisited this concept and provided a more accessible explanation. According to this model, any source of noise generates a packet of leading spiral waves, which are then amplified as they swing from leading to trailing. The amplified trailing waves propagate through the disk center and reemerge as a leading disturbance, which then triggers the next swing amplification loop. The repeated amplification process amplifies minor noise into a prominent spiral pattern and ultimately facilitates the formation of a coherent bar. The positive feedback loop predicts exponential growth of the spiral/bar strength.
@@
Several studies investigated the growth rate or equivalently the formation timescale of bars formed via internal instability. The formation timescale is found to depend on the bulge-to-disc mass fraction (Kataria & Das, 2018), the Toomre $Q$ parameter (Hozumi, 2022; Worrakitpoonpon, 2024), and the disk thickness (Ghosh et al., 2023). Fujii et al. (2018) found an exponentially decreasing relation between the bar formation timescale and decreasing disc mass fraction, which was then confirmed in Bland-Hawthorn et al. (2023). An important improvement in the latter work is the quantification of the exponential growth timescale associated with the swing amplification loop instead of Fujii’s fixed threshold for when a bar appears ($A_{2,max}>0.2$).
@@
Besides the internal instability, bar formation can also be externally induced by various gravitational perturbations, including flyby encounters with other galaxies, collisions, mergers, or the tidal effects of galaxy clusters (Byrd et al., 1986; Gauthier et al., 2006; Martinez-Valpuesta et al., 2017; Łokas et al., 2016; Łokas, 2019b). A comprehensive investigation is desired to compare the growth and evolution of bars formed through different mechanisms. Studies have indicated that tidal interactions can either promote or delay bar formation (Łokas et al., 2016; Moetazedian et al., 2017; Łokas, 2018; Pettitt & Wadsley, 2018; Zana et al., 2018). This raises a question: if tidal forces either promote or delay the formation of bars, do they result in bars that grow at a faster or slower rate compared to those formed internally? Alternatively, do these bars grow at similar rates but onset at different times? Moetazedian et al. (2017) found that the exponential growth rate of tidal bars is independent of the number of satellite encounters and their orbits. However, the rescaled mass of the primary satellite in their simulation is only $<1\;\%$ compared to the host halo mass of the galaxy, which limits the strength of the tidal perturbation.
@@
Previous studies have predominantly focused on either the internal or external mechanisms of bar formation, with less attention given to comparing these two mechanisms. In this series of papers, we aim to offer a comprehensive comparison between bars formed through different mechanisms. In Zheng & Shen (2024, Paper I hereafter), we conducted a series of controlled $N$-body simulations to examine the pattern speed $\Omega_{p}$ of tidally induced bars and compared them with their spontaneous counterparts in the same galaxy model. We found that bars formed via both mechanisms exhibit similar distributions in the pattern speed–bar strength ($\Omega_{p}-A_{2}$) space and display comparable ratios of co-rotation radius to bar length (${\cal R}={R_{\mathrm{CR}}}/{R_{\mathrm{bar}}}$). When “tidally induced bars” refer specifically to those in galaxies stable against bar instability, they indeed rotate slower than internally developed ones in less stable galaxies. However, this difference in pattern speed is attributed to the varying internal properties of the bar host galaxies, rather than the different formation mechanisms. In this second paper, we broaden our investigation to compare the growth timescales and onset times between tidally induced and spontaneously developed bars. Our objective is to understand how external perturbations influence the growth of bar structures.
@@ 2 Simulations
We calculate the Toomre $Q$ parameter to evaluate the stability of the disks and observe similar $Q$ curves across all three disk models. The cold, warm, and hot disks exhibit minimum Toomre stability values ($Q_{\mathrm{min}}$) of 0.82, 1.34, and 2.24, respectively. According to the criterion for bar formation in Jang & Kim (2023), the cold disk is susceptible to bar instability, the warm disk is marginally stable, and the hot disk is stable against bar formation. In isolated simulations, the cold and warm disks are capable of spontaneously forming bars, with the warm disk requiring more time to reach a comparable bar strength. The hot disk does not develop a bar within 6 Gyr.
Utilizing these galaxy models, we perform a series of flyby interaction simulations. The perturber in these simulations is a pure Hernquist DM halo composed of live particles. The galaxy and perturber are placed on a hyperbolic orbit, inspired by the configuration presented by Łokas (2018) (refer to her Figure 1 for a schematic). We investigate mass ratios of the perturber to the galaxy ($M_{\mathrm{per}}/M_{\mathrm{gal}}$) of 1/1, 1/3, and 1/10, resulting in strong, moderate, and weak tidal perturbations, respectively. For each mass ratio, we adjust the inclination angle between the stellar disk plane and the perturber’s orbit plane to simulate prograde ($0^{\circ}$), perpendicular ($90^{\circ}$), and retrograde ($180^{\circ}$) flyby interactions. The closest approach is set to occur at approximately $t_{\rm peri}=0.5$ Gyr for all flyby simulations.
Each simulation is evolved for 6 Gyr with GADGET-4 (Springel, 2005; Springel et al., 2021). We populate the stellar disk with $0.5$ million particles and the galaxy DM halo with $1$ million particles, leading to particle masses of $7.2\times 10^{4}\;{\rm M}_{\odot}$ for the disk and $3.6\times 10^{5}\;{\rm M}_{\odot}$ for the halo. The gravitational softening lengths are set at 23 pc for stars and 57 pc for DM particles.
@@ 3 Bar growth timescale
Figure 1: Exponential fitting to the $A_{2}$ growth. The black solid lines represent the evolution of $A_{2}$, with the $1\sigma$ range highlighted by the grey area. The blue solid lines indicate the results of the exponential fitting. The regions selected for fitting are denoted by green vertical dotted lines, while the intended pericenter time, $t_{\rm peri}=0.5$ Gyr, is marked by red vertical dashed lines. The inset axes display the $A_{2}$ evolution across the entire simulation duration of 6 Gyr. The text box notes the simulation information of mass ratio ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/10$, 1/3, and 1/1) and orbit type (prograde, perpendicular, and retrograde).
@@
where $m_{i}$ and $\phi_{i}$ represent the mass and the azimuthal angle of the $i$-th particle, respectively. The summation is conducted over the stellar particles within a cylindrical region where $R\leq 4\;R_{d}$, i.e. $R\leq 8\;{\rm kpc}$ in our galaxy models. The bar strength $A_{2}$ is then calculated as:
@@
We fit the $A_{2}$ evolution with an exponential function:
@@
where $C=0.1\exp(-t_{0}/\tau_{\mathrm{bar}})$. We prefer Equation 3 for fitting the $A_{2}$ evolution due to its more intuitive interpretation of the fitted parameters. We explored other fitting functions in Appendix A. These different functions yield similar $\tau_{\mathrm{bar}}$ values with differences of approximately 20% while maintaining equivalent fitting quality.
We perform the fitting with the Levenberg–Marquardt algorithm through the curve_fit function in the scipy.optimize package (Virtanen et al., 2020). The fitting range starts when the bar strength reaches $A_{2}=0.025$ and ends when $A_{2}$ reaches $0.8\;A_{2,peak}$ with $A_{2,peak}$ being the peak $A_{2}$ value during the bar formation stage. This lower limit is chosen to avoid initial stages when $A_{2}$ cannot be well decided. The upper limit is set to avoid the saturation of the bar strength at which the bar growth deviates from the exponential profile. We manually introduce a constant error [^1] The exact value of $\sigma_{A2}$ is not crucial as long as it is reasonably chosen to avoid unrealistic impact on the fitting results. of $\sigma_{A2}=0.04$ to the $A_{2}$ values to account for noise and to calculate the $1\sigma$ error of the fitting parameters.
Our fitting results are presented in Figure 1. The majority of the simulations show nice exponential growth profiles of the bar strength with the fitting results well overlapping the $A_{2}$ evolution, consistent with the swing amplifier feedback loop theory. However, several simulations show a deviation from this exponential growth, predominantly occurring in prograde interactions with massive perturbers. Visual examination of the stellar surface density maps reveals that this deviation is primarily due to the substantial contribution of spiral arms to the $A_{2}$ signal. We further validate the reliability of our fitting method by comparing the fitted $t_{0}$ values against the true $t_{0}$ in Appendix A. This comparison demonstrates a strong agreement between the fitted and true $t_{0}$ values, with most simulations exhibiting an offset of less than 0.1 Gyr.
@@ 4.1 Do tidal bars grow faster?
Figure 2: Results of the exponential fitting for the cold disk model (left column) and the warm disk model (right column). Each scatter point represents a simulation and dashed lines indicate the results for the same disk model in isolation. Top panels: the growth timescale $\tau_{\mathrm{bar}}$. Tidally induced bars exhibit $\tau_{\mathrm{bar}}$ values similar to those of internally developed bars, with an offset of less than 20%, suggesting that tidal bars and spontaneous bars in the same disk have comparable growth rates. Middle panels: the fitted $t_{0}$, representing the time at which $A_{2}$ reaches 0.1. Tidal perturbations promote bar formation by advancing the onset time, as evidenced by their smaller $t_{0}$ values. Bottom panels: same as the middle panels, but for the true $t_{0}$.
Figure 3: The $A_{2}$ evolution against $(t-t_{0,\ fit})$. All tidally induced bars, except those in prograde interactions (red lines) with massive perturbers where spiral arms significantly contribute to the $A_{2}$ signal, exhibit growth patterns very similar to the isolated case. This similarity confirms that tidal bars share comparable growth rates with their spontaneous counterparts in the same disk.
@@
The comparison is presented in Figure 2 with the left column for the cold disk and the right column for the warm disk. The upper panels show that the tidally induced bars exhibit $\tau_{\mathrm{bar}}$ comparable to that of the internally developed bars in the same disk. In most cases, the discrepancy is less than 20%. Considering that different fitting functions can yield $\tau_{\mathrm{bar}}$ differences of approximately 20%, our result suggests that tidal bars share similar growth rates with their spontaneous counterpart in the same disk. Additionally, the relatively small $1\sigma$ error bar for $\tau_{\mathrm{bar}}$ indicates that the fitting results are robust against noise in the $A_{2}$ evolution.
@@
The largest discrepancy in $\tau_{\mathrm{bar}}$ is observed in prograde interactions with a mass ratio of 1/1. In such cases, the perturbation is sufficiently strong to generate prominent spiral arms that significantly contribute to the $A_{2}$ signal. This situation reduces the reliability of the fitted $\tau_{\mathrm{bar}}$ as it no longer exclusively represents the growth of the bar structure. It is beyond the scope of this study to disentangle the contribution of the spiral arms to the $A_{2}$ signal from the bar growth. Therefore, we conclude that tidally induced bars grow as fast as their spontaneous counterpart in the same disk while acknowledging the potential influence of spiral arms on the $A_{2}$ signal.
@@
In the middle panels of Figure 2, we present the fitted $t_{0}$, the time at which the fitted $A_{2}$ profile reaches 0.1. Tidally induced bars have an earlier onset of formation, as indicated by their smaller $t_{0}$ values. The onset of bar formation occurs earlier with stronger tidal perturbations. The only exception happens in the perpendicular interaction of cold disk with perturber of $M_{\mathrm{per}}/M_{\mathrm{gal}}$$=1/3$, where bar formation is delayed by $\sim 0.4$ Gyr. The underlying cause of the delay will be explored in subsection 4.2.
@@
In the cold disk model (middle left panel), the prograde interactions exhibit significantly smaller $t_{0}$ values compared to the isolated case, whereas the perpendicular and retrograde interactions show similar $t_{0}$ values. These results indicate that bar formation in the cold disk is predominantly driven by internal disk instability, with only severe external perturbations capable of accelerating the emergence of bars.
@@
In the warm disk model (middle right panel), all interactions result in smaller $t_{0}$ values compared to the isolated case except for the weakest interactions with $M_{\mathrm{per}}/M_{\mathrm{gal}}=1/10$ on perpendicular and retrograde orbits. The warm disk is only marginally stable against bar formation in isolation. In its interaction simulations, bar formation is a product of the combination of both internal disk instability and external perturbations.
In the lower panels of Figure 2, we show the true $t_{0}$, the time when the actual $A_{2}$ profile reaches 0.1. These results corroborate the findings from the fitted $t_{0}$.
@@
Figure 2 illustrates that tidal interactions promote/delay bar formation by either advancing/postponing the onset of this process. Following the onset of bar formation, tidally induced bars grow at a rate comparable to internally developed bars within the same disk. This conclusion is further reinforced by Figure 3, where we plot the $A_{2}$ evolution against $(t-t_{0,\mathrm{fit}})$. All tidally induced bars exhibit similar growth patterns when referenced against the isolated case, regardless of whether the tidal perturbation accelerates or postpones bar formation. The only notable deviation occurs in prograde interactions with massive perturbers, where the contribution of spiral arms to $A_{2}$ is significant. In such scenarios, the tidal force exerted by the perturber impacts stars in the prograde disk for a longer duration than those in the perpendicular or retrograde disk, thereby promoting the development of spiral arms. We refer the reader to Łokas (2018) and Paper I for more discussion on the effect of the inclination of the perturber orbit.
@@
The consistency in $\tau_{\mathrm{bar}}$ suggests that the growth of tidal bars is predominantly governed by the inherent nature of the galaxy, which remains largely unaffected by the tidal perturbation. Tidal perturbations do not substantially alter the swing amplification process, although they provide stronger initial noises that trigger the swing amplification loop. Tidally induced bars grow as fast as their spontaneous counterpart in the same disk.

### v3

# Comparison of bar formation mechanisms. II. does a tidally induced bar grow faster than an internally developed bar?

@@ Abstract
Bar structures can form internally due to the instability of their host galaxies or externally due to perturbations from other galaxies. We systematically quantify the growth timescales ($\tau_{\mathrm{bar}}$) of bars formed through these two mechanisms with a series of controlled $N$-body simulations. In galaxies susceptible to bar instability, tidally induced bars display $\tau_{\mathrm{bar}}$ values comparable to those of internally developed bars within the same disk. Tidal perturbations promote(delay) bar formation by advancing(postponing) its onset, but the growth rate of the bar structure remains largely unchanged. In these interaction scenarios, the bar formation is still driven primarily by the galaxy’s internal nature, which remains unaffected by tidal perturbations. As the external perturbation wave reaches the galaxy’s center, it evokes a “seed bar” that is then swing amplified. In this scenario, the onset of bar formation is advanced. Conversely, bar formation may be delayed if the external perturbation wave is out of phase with the preexisting spontaneously developed seed bar, which causes destructive interference and limits the bar growth. In the hot disk model that resists bar formation in isolation, the $\tau_{\mathrm{bar}}$ of the tidally forced bar correlates with the strength of the perturbation. The bar growth in this model deviates from an exponential profile and is better described by a linear function. The varied $\tau_{\mathrm{bar}}$ and the preference for linear growth contrast with bars formed in galaxies inherently susceptible to bar instability. These tidally forced bars may not adhere to the swing amplification mechanism that predicts an exponential bar growth.
@@ 1 Introduction
Bars are a common feature in spiral galaxies. Locally, about half of spiral galaxies possess bars in optical bands, with this proportion rising to two-thirds in infrared bands (Marinova & Jogee, 2007; Menéndez-Delmestre et al., 2007; Erwin, 2018; Lee et al., 2019). Although a smaller fraction of galaxies have bars at higher redshifts, they still represent 13% in the range $2<z\leq 3$ (Le Conte et al., 2024) and appear as early as $z\sim 4$ (Guo et al., 2024). Bars significantly influence the evolution of their host galaxies, including triggering gas inflow, affecting star formation, and contributing to the formation of pseudobulges (Masters et al., 2011; Li et al., 2015; Lin et al., 2017, 2020; Iles et al., 2022). The Milky Way also has a bar (e.g., de Vaucouleurs, 1964; Blitz & Spergel, 1991), and its boxy bulge and gas dynamics are directly linked to this bar (e.g., Shen et al., 2010; Li et al., 2018). The prevalence and impact of bars on galaxies highlight the importance of studying their formation and evolution.
@@
Galactic bars can form via internal or external mechanisms. Internally, bars can develop spontaneously due to the gravitational instability of the disks (Hohl, 1971; Ostriker & Peebles, 1973; Sellwood, 2014; Łokas, 2019a). A widely recognized mechanism for spontaneous bar formation is the swing amplification loop, initially developed by Goldreich & Lynden-Bell (1965) and Julian & Toomre (1966) and further elaborated by Toomre (1981). Binney (2020) revisited this concept and provided a more accessible explanation. According to this model, any source of noise generates a packet of leading spiral waves, which are then amplified as they swing from leading to trailing. The amplified trailing waves propagate through the disk center and reemerge as a leading disturbance, which then triggers the next swing amplification loop. The repeated amplification process amplifies minor noise into a prominent spiral pattern and ultimately facilitates the formation of a coherent bar. The positive feedback loop predicts exponential growth of the spiral/bar strength.
@@
Several studies investigated the growth rate or equivalently the formation timescale of bars formed via internal instability. The formation timescale is found to depend on the bulge-to-disk mass fraction (Kataria & Das, 2018), the Toomre $Q$ parameter (Hozumi, 2022; Worrakitpoonpon, 2024), and the disk thickness (Ghosh et al., 2023). Fujii et al. (2018) found an exponentially decreasing relation between the bar formation timescale and increasing disk mass fraction. The “Fujii relation” was then confirmed in Bland-Hawthorn et al. (2023). An important improvement in the latter work is the quantification of the exponential growth timescale associated with the swing amplification loop instead of Fujii’s fixed threshold for when a bar appears ($A_{2,max}>0.2$). Bland-Hawthorn et al. (2023) demonstrated that an exponential fit to the growth of $A_{2}$ over time provides a timescale and an onset time, and also reproduces the Fujii relation independently of any assumptions about $A_{2}$. Several studies have suggested that the disk mass fraction plays a primary role in determining the bar formation timescale (Fujii et al., 2018; Bland-Hawthorn et al., 2023, 2024, 2025). Thus, it is important to control the disk mass fraction when investigating the effects of other parameters on bar formation.
@@
Besides the internal instability, bar formation can also be externally induced by various gravitational perturbations, including flyby encounters with other galaxies, collisions, mergers, or the tidal effects of galaxy clusters (Byrd et al., 1986; Gauthier et al., 2006; Łokas et al., 2016; Martinez-Valpuesta et al., 2017; Łokas, 2019b). A comprehensive investigation is desired to compare the growth and evolution of bars formed through different mechanisms. Studies have indicated that tidal interactions can either promote or delay bar formation (Łokas et al., 2016; Moetazedian et al., 2017; Łokas, 2018; Pettitt & Wadsley, 2018; Zana et al., 2018). This raises a question: if tidal forces either promote or delay the formation of bars, do they result in bars that grow at a faster or slower rate compared to those formed internally? Alternatively, do these bars grow at similar rates but onset at different times? Moetazedian et al. (2017) found that the exponential growth rate of tidal bars is independent of the number of satellite encounters and their orbits. However, the rescaled mass of the primary satellite in their simulation is only $<$1% compared to the host halo mass of the galaxy, which limits the strength of the tidal perturbation.
@@
Previous studies have predominantly focused on either the internal or external mechanisms of bar formation, with less attention given to comparing these two mechanisms. In this series of papers, we aim to offer a comprehensive comparison between bars formed through different mechanisms. In Zheng & Shen (2025, Paper I hereafter), we conducted a series of controlled $N$-body simulations to examine the pattern speed $\Omega_{p}$ of tidally induced bars and compared them with their spontaneous counterparts in the same galaxy model. We found that bars formed via both mechanisms exhibit similar distributions in the pattern speed–bar strength ($\Omega_{p}-A_{2}$) space and display comparable ratios of corotation radius to bar length (${\cal R}={R_{\mathrm{CR}}}/{R_{\mathrm{bar}}}$). As for “tidally forced bars” that refer specifically to those in galaxies stable against bar instability, they indeed rotate slower than internally developed ones in less stable galaxies. However, this difference in pattern speed is attributed to the varying internal properties of bar host galaxies, rather than the different formation mechanisms. In this second paper, we broaden our investigation to compare the growth timescales and onset times between tidally induced and spontaneously developed bars. Our objective is to understand how external perturbations influence the growth of bar structures.
@@ 2 Simulations
Following Fujii et al. (2018) and Bland-Hawthorn et al. (2023), we calculate the $f_{\text{disk}}$ parameter as the ratio of the disk mass to the total galaxy mass within the radius at $R=2.2\;R_{d}$ at which the rotation curve roughly peaks. $f_{\text{disk}}$ is defined by
$f_{\text{disk}}=\frac{V^{2}_{\text{c,disk}}(R)}{V^{2}_{\text{c,tot}}(R)}\Bigg{|}_{R=2.2\;R_{d}},$ (1)
with $V_{\text{c}}(R)$ denoting the circular velocity at radius $R$. The cold, warm, and hot disks have $f_{\text{disk}}$ values of 0.47, 0.47, and 0.44, respectively, positioning them within the intermediate disk mass fraction range on the Fujii relation (Fujii et al., 2018; Bland-Hawthorn et al., 2023).
These models have very similar $f_{\text{disk}}$ values, which is an expected outcome of using identical parameters for the stellar disk and DM halo density profiles. The hot disk model has a slightly lower $f_{\text{disk}}$ value due to its slightly reduced circular speed in the inner region. The rationale behind the hot disk model’s lower circular speed is elaborated in Section 2.1 and Figure 1 of Paper I.
With $f_{\text{disk}}$ being very similar among the models, the Toomre $Q$ parameter is important in evaluating the stability of the disks. The $Q$ curves show a similar shape across all three disk models but hold different values. The cold, warm, and hot disks exhibit minimum Toomre stability values ($Q_{\mathrm{min}}$) of 0.82, 1.34, and 2.24, respectively. According to the criterion for bar formation in Jang & Kim (2023), the cold disk is susceptible to bar instability, the warm disk is marginally stable, and the hot disk is stable against bar formation. In isolated simulations, the cold and warm disks are capable of spontaneously forming bars, with the warm disk requiring more time to reach a comparable bar strength. The hot disk does not develop a bar within 6 Gyr.
Utilizing these galaxy models, we perform a series of flyby interaction simulations. The perturber in these simulations is a pure Hernquist DM halo composed of live particles. The galaxy and perturber are placed on a hyperbolic orbit, inspired by the configuration presented by (Łokas, 2018, refer to her Figure 1 for a schematic) . We investigate mass ratios of the perturber to the galaxy ($M_{\mathrm{per}}/M_{\mathrm{gal}}$) of 1/1, 1/3, and 1/10, resulting in strong, moderate, and weak tidal perturbations, respectively. For each mass ratio, we adjust the inclination angle between the stellar disk plane and the perturber’s orbit plane to simulate prograde ($0^{\circ}$), perpendicular ($90^{\circ}$), and retrograde ($180^{\circ}$) flyby interactions. The closest approach is set to occur at approximately $t_{\rm peri}=0.5$ Gyr for all flyby simulations.
Each simulation is evolved for 6 Gyr with GADGET-4 (Springel, 2005; Springel et al., 2021). We populate the stellar disk with $0.5$ million particles and the galaxy DM halo with one million particles, leading to particle masses of $7.2\times 10^{4}\;{\rm M}_{\odot}$ for the disk and $3.6\times 10^{5}\;{\rm M}_{\odot}$ for the halo. The gravitational softening lengths are set at 23 pc for stars and 57 pc for DM particles. We acknowledge that the mass resolution of our simulations is not exceptionally high, yet it is adequate to capture the growth of the bar. Dubinski et al. (2009) demonstrated that the evolution of the bar exhibits convergent behavior when the number of halo particles ranges from one million to 10 million. This convergence is evident when studying bar growth, pattern speed evolution, the dark matter halo density profile, and the nonlinear analysis of orbital resonances.
@@ 3 Bar growth timescale
Figure 1: Exponential fitting to the $A_{2}$ growth. The black solid lines represent the evolution of $A_{2}$, with the $1\sigma$ range highlighted by the gray area. The blue solid lines indicate the results of the exponential fitting. The regions selected for fitting are denoted by green vertical dotted lines, while the intended pericenter time, $t_{\rm peri}=0.5$ Gyr, is marked by red vertical dashed lines. The inset axes display the $A_{2}$ evolution across the entire simulation duration of 6 Gyr. The text box notes the simulation information of mass ratio ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/10$, 1/3, and 1/1) and orbit type (prograde, perpendicular, and retrograde).
@@
where $m_{i}$ and $\phi_{i}$ represent the mass and the azimuthal angle of the $i$th particle, respectively. The summation is conducted over the stellar particles within a cylindrical region where $R\leq 4\;R_{d}$, i.e., $R\leq 8\;{\rm kpc}$ in our galaxy models. The bar strength $A_{2}$ is then calculated as
@@
Following Bland-Hawthorn et al. (2023), we fit the $A_{2}$ evolution with an exponential function:
@@
where $C=0.1\exp(-t_{0}/\tau_{\mathrm{bar}})$. When applying to the same data set, both fitting functions yield the same value and the same standard deviation error of the growth timescale $\tau_{\mathrm{bar}}$. Neither function is better than the other, but we prefer Equation 4 for fitting the $A_{2}$ evolution due to the slightly better physical interpretability of $t_{0}$ compared to $C$. Besides, the $t_{0}$ parameter assists in indicating the initiation time of bar formation, which is beneficial when comparing bar formation scenarios with similar $\tau_{\mathrm{bar}}$ values. This is helpful when comparing tidally induced bars with the bars formed in isolated simulations within the same disk model as discussed in section 4. We explored other fitting functions in the Appendix A. These different functions yield similar $\tau_{\mathrm{bar}}$ values with differences of approximately 20% while maintaining equivalent fitting quality.
We perform the fitting with the Levenberg–Marquardt algorithm through the curve_fit function in the scipy.optimize package (Virtanen et al., 2020) , following the same method as Bland-Hawthorn et al. (2023). The fitting range starts when the bar strength reaches $A_{2}=0.025$ and ends when $A_{2}$ reaches $0.8\;A_{2,peak}$ with $A_{2,peak}$ being the peak $A_{2}$ value during the bar formation stage.  [^1] Only the bar formation stage before the bar buckling phase is considered for the fitting. The secular growth stage after buckling is excluded. This lower limit is chosen to avoid initial stages when $A_{2}$ cannot be well decided. The upper limit is set to avoid the saturation of the bar strength at which the bar growth deviates from the exponential profile. We manually introduce a constant error [^2] The exact value of $\sigma_{A2}$ is not crucial as long as it is reasonably chosen to avoid unrealistic impact on the fitting results. of $\sigma_{A2}=0.04$ to the $A_{2}$ values to account for noise and to calculate the $1\sigma$ error of the fitting parameters.
Our fitting results are presented in Figure 1. The majority of the simulations show exponential growth profiles of the bar strength with the fitting results well overlapping the $A_{2}$ evolution, consistent with the swing amplifier feedback loop theory. However, several simulations show a deviation from this exponential growth, predominantly occurring in prograde interactions with massive perturbers. Visual examination of the stellar surface density maps reveals that this deviation is primarily due to the substantial contribution of spiral arms to the $A_{2}$ signal. We provide more details regarding our fitting method and further demonstrate its reliability in the Appendix A.
@@ 4.1 Do tidal bars grow faster?
Figure 2: Results of the exponential fitting for the cold disk model (left column) and the warm disk model (right column). Each scatter point represents a simulation and the error bar shows the standard deviation on the parameter estimate. The dashed lines indicate the results for the same disk model in isolation. Upper panels: the growth timescale $\tau_{\mathrm{bar}}$. Tidally induced bars exhibit $\tau_{\mathrm{bar}}$ values similar to those of internally developed bars, with an offset of less than 20%, suggesting that tidal bars and spontaneous bars in the same disk have comparable growth rates. Lower panels: the fitted $t_{0}$, representing the time at which $A_{2}$ reaches 0.1. Tidal perturbations promote bar formation by advancing the onset time, as evidenced by their smaller $t_{0}$ values.
Figure 3: The $A_{2}$ evolution against $(t-t_{0})$. All tidally induced bars, except those in prograde interactions (red lines) with massive perturbers where spiral arms significantly contribute to the $A_{2}$ signal, exhibit growth patterns very similar to the isolated case. This similarity confirms that tidal bars share comparable growth rates with their spontaneous counterparts in the same disk.
@@
The comparison is presented in Figure 2 with the left column for the cold disk and the right column for the warm disk. The upper panels show that the tidally induced bars exhibit $\tau_{\mathrm{bar}}$ comparable to that of the internally developed bars in the same disk. In most cases, the discrepancy is less than 20%. Considering that different fitting functions can yield $\tau_{\mathrm{bar}}$ differences of approximately 20%, our result suggests that tidal bars share similar growth rates with their spontaneous counterpart in the same disk. Additionally, the relatively small standard deviation of $\tau_{\mathrm{bar}}$ indicates that the fitting results are robust against noise in the $A_{2}$ evolution.
@@
The largest discrepancy in $\tau_{\mathrm{bar}}$ is observed in prograde interactions with a mass ratio of 1/1. In such cases, the perturbation is sufficiently strong to generate prominent spiral arms that significantly contribute to the $A_{2}$ signal. This situation reduces the reliability of the fitted $\tau_{\mathrm{bar}}$ as it no longer exclusively represents the growth of the bar structure. Bland-Hawthorn (2025, in preparation) reports bars and spiral arms may be separated by radial actions. However, it is beyond the scope of this study to disentangle the contribution of the spiral arms to the $A_{2}$ signal from the bar growth. Therefore, we conclude that tidally induced bars grow as fast as their spontaneous counterpart in the same disk while acknowledging the potential influence of spiral arms on the $A_{2}$ signal.
@@
In the lower panels of Figure 2, we present the fitted $t_{0}$, the time at which the fitted $A_{2}$ profile reaches 0.1. Tidally induced bars have an earlier onset of formation, as indicated by their smaller $t_{0}$ values. The onset of bar formation occurs earlier with stronger tidal perturbations. The only exception happens in the perpendicular interaction of the cold disk with a perturber of $M_{\mathrm{per}}/M_{\mathrm{gal}}$$=1/3$, where bar formation is delayed by $\sim 0.4$ Gyr. The underlying cause of the delay will be explored in subsection 4.2.
@@
In the cold disk model (lower left panel), the prograde interactions exhibit significantly smaller $t_{0}$ values compared to the isolated case, whereas the perpendicular and retrograde interactions show similar $t_{0}$ values. These results indicate that bar formation in the cold disk is predominantly driven by internal disk instability, with only severe external perturbations capable of accelerating the emergence of bars.
@@
In the warm disk model (lower right panel), all interactions result in smaller $t_{0}$ values compared to the isolated case except for the weakest interactions with $M_{\mathrm{per}}/M_{\mathrm{gal}}=1/10$ on perpendicular and retrograde orbits. The warm disk is only marginally stable against bar formation in isolation. In its interaction simulations, bar formation is a product of the combination of both internal disk instability and external perturbations.
@@
Figure 2 illustrates that tidal interactions promote(delay) bar formation by either advancing(postponing) the onset of this process. Following the onset of bar formation, tidally induced bars grow at a rate comparable to internally developed bars within the same disk. This conclusion is further reinforced by Figure 3, where we plot the $A_{2}$ evolution against $(t-t_{0})$. All tidally induced bars exhibit similar growth patterns when referenced against the isolated case, regardless of whether the tidal perturbation accelerates or postpones bar formation. The only notable deviation occurs in prograde interactions with massive perturbers, where the contribution of spiral arms to $A_{2}$ is significant. In such scenarios, the tidal force exerted by the perturber impacts stars in the prograde disk for a longer duration than those in the perpendicular or retrograde disk, thereby promoting the development of spiral arms. We refer the reader to Łokas (2018) and Paper I for more discussion on the effect of the inclination of the perturber’s orbit.
@@
The consistency in $\tau_{\mathrm{bar}}$ suggests that the growth of tidal bars is predominantly governed by the inherent nature of the galaxy, which remains largely unaffected by the tidal perturbation. Tidal perturbations do not substantially alter the swing amplification process, although they provide stronger initial noise that triggers the swing amplification loop. Tidally induced bars grow as fast as their spontaneous counterpart in the same disk.